OpenStudy (anonymous):

In a triangle ABC, AD is the bisector of angle BAC and angle C is greater than angle B. Prove that angle ADB > angle ADC.

OpenStudy (anonymous):



OpenStudy (anonymous):

Please help I have an exam tomorrow and need to get my doubts cleared.

OpenStudy (anonymous):

@lgbasallote please please help ....

OpenStudy (anonymous):

@lgbasallote help!!!!!

OpenStudy (anonymous):

Can you help @Mimi_x3?

OpenStudy (wwe123):

one min

OpenStudy (wwe123):

AD is the bisector of angle BAC , so angle BAD = angle CAD angle C >angle B angle ADB =180-(ABD+BAD) angle ADC = 180-(CAD+ACD) since angle BAD = angle CAD angle ADB =180-(BAD+ABD) angle ADC =180-(BAD+ACD) angle C >angle B hence, angle ADB > angle ADC.

OpenStudy (anonymous):

Thanks wwe123.